OpenStudy (anonymous):

Find the limit as x approaches 3?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3} = (\sqrt{2x+3}-\sqrt{3x})/(x^2-3x)\]

OpenStudy (anonymous):

I know the limit is -0.0555 or -1/18 from graphing it but I can't figure out how to find the limit algebraically

jimthompson5910 (jim_thompson5910):

Try multiplying top and bottom by \(\Large \sqrt{2x+3}+\sqrt{3x}\) to rationalize the numerator.

OpenStudy (anonymous):

okay thank you! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me know if you're able to finish up or not

OpenStudy (anonymous):

i was able to finish it up, thanks again!

jimthompson5910 (jim_thompson5910):

you're welcome